All three ex81 zeros were reset. This was done between 1-2pm.

We have been seeing non-zero flow rates when there is no flow.

Currently, Center bay seems to be metering less flow with quick initial test with MP2000.

Further evaluation will be done.

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 13.10.5 - Documentation